10 typist in room give rise to an average sound intensity of 58db, what is the intensity level in the room when an additional three typist each generating the same amount of noise being to type?
1
Expert's answer
2018-02-05T08:15:08-0500
L=58 dB=10 log⁡〖10I/I_0 〗
L^'=10 log⁡〖13I/I_0 〗=10(log⁡〖10I/I_0 〗+log⁡1.3 )=58+10 log⁡1.3=59.1 dB.